the normalized wave function for the first excited state of particle in one dimensional box


1
Expert's answer
2021-04-08T07:35:47-0400

The wave function can be written as

"\Psi_n(x)=A_nsin(\dfrac{n\pi x}{L})"


The constant "A_n" is determined by normalization condition

"\int\Psi^2_n dx=\int A_n^2sin^2(\dfrac{n\pi x}{L})dx=1"


The result of evaluating the integral and solving for "A_n=\sqrt{\dfrac{2}{L}}" is independent from n.


The normalized wave function for a particle in a box are thus

"\Psi_n(x)=\sqrt{\dfrac{2}{L}}sin(\dfrac{n\pi x}{L})"

For first excited state n = 2

"\therefore\space \Psi_2(x)=\sqrt{\dfrac{2}{L}}sin(\dfrac{2\pi x}{L})"